OpenStudy (anonymous):

If tan theta= 15/8, then: A. sec theta = 17/8 B. cos theta = 15/17 C. cot theta = 8/15 D. csc theta = 17/15 There's potentially more than one answer.

OpenStudy (johnweldon1993):

Well we know \[\large tan = \frac{sin}{cos} = \frac{15}{8}\] This tells us that sin = 15 and cos = 8

OpenStudy (johnweldon1993):

So a good hint, cos is not represented as a fraction..so B does not make sense

OpenStudy (johnweldon1993):

Now, Lets begin with A \[\large sec \theta = \frac{1}{cos} = \frac{1}{8}\] Secant can be written as 1/cos...and since we know cos = 8 we would have 1/8 So A is out as well

OpenStudy (johnweldon1993):

Now D \[\large csc\theta = \frac{1}{sin} = \frac{1}{15}\] cosecant can be written as 1/sin and since we know sin = 15... So D is out

OpenStudy (johnweldon1993):

And finally C \[\large \cot\theta = \frac{1}{tan} = \frac{1}{\frac{sin}{cos}} = \frac{cos}{sin} = \frac{8}{15}\] Here we have our answer!

OpenStudy (anonymous):

Is the key just memorizing when to use sin/cos/tan and SOACAHTOA or whatever?

myininaya (myininaya):

it is incorrect to say sin theta is 15 and cos of theta is 8. the outputs should be between -1 and 1

myininaya (myininaya):

build a right triangle base on what is given then find the hyp

myininaya (myininaya):

then you can find sin of theta cos of theta and so on based off that triangle

OpenStudy (anonymous):

Refer to the one line calculation using Mathematica v9.